Consider a reversible heat engine that employs a hot reservoir at a temperature of 660 K and a cold reservoir at 230 K.

(a) What is the entropy change of the hot reservoir during a period in which 4300 J is extracted from the hot reservoir?

(b) What is the change in the entropy of the cold reservoir?

(c) Find the change in the entropy of the engine itself during this time.
Answer #1

a ) s = -Q/T = 4870 /660 = - 7.37878788

b) + 7.37878788 ( Because this is a REVERSIBLE heat engine, there can be no entropy generation in the engine. All entropy in = all entropy out. )

c ) Zero. Because this is a REVERSIBLE heat engine, there can be no entropy generation in the engine. All entropy in = all entropy out.